OpenStudy (anonymous):

Can someone find the indefinate integral of (dy/dx)= x (x -5)^1/2?

OpenStudy (anonymous):

\[(2/15) (-5+x)^{3/2}(10+3x)\]

OpenStudy (anonymous):

Using substitution... \[\Large \begin{array}{l} \int {x{{(x - 5)}^{1/2}}} dx\\ u = (x - 5)\\ u + 5 = x\\ \frac{{du}}{{dx}} = 1\\ du = 1dx = dx\\ \int {(u + 5)} \cdot {u^{1/2}}du\\ = \int {{u^{3/2}} + 5{u^{1/2}}du} \\ = \int {{u^{3/2}}du} + \int {5{u^{1/2}}} du\\ = \frac{{{u^{5/2}}}}{{5/2}} + \frac{{5{u^{3/2}}}}{{3/2}} + c\\ = \frac{{2{u^{5/2}}}}{5} + \frac{{10{u^{3/2}}}}{3} + c \end{array}\] Substitute back in for u and simplify.